2009 October 23 LBC Tech Night (active optics work - see PLAN)

Times are UT:

00:00 Took a series of 25 biases with Blue/Red - these appear in newdata and in Repository/20091023. Lights were off.

01:17 Slew to pointing star to focus/collimate and check pointing. ACT0287. Initial CA and IE are -46.9 and -10.3. Red filter wheel #2 more than 10 times from position - 2nd attempt it moved to position OK.

B no good pupils found - too light still. R sent +8817 Z4. Lot of coma too. Still blue is like a bias.... taking the Pointing Start 014241 (B) and 014240 (R) to check. Blue image devoid of stars, even the bright pointing star. Took 30 sec images at V and r: 014655 and 014652. Blue image is devoid of stars and has mean of about 700 adu - like a dark or bias. Shutter?

02:00 Power cycled Other systems and HK --- try again: R 020033 and B 020036. Now Blue is good.

02:04 Focus/Collimate now... manual correction needed (020451 has coma and defocus and FPIA fit is poor. Give Z4=+6000nm). Backout... 021016 and 021013.

Gave Z7 = -1500nm to both SX and DX... repeat

FPIA crashes on Blue (021832) - gave another -1000nm Z7. New CA = -61.5, New IE = 3.4. Not too far off.

In last DOFPIA run, +100 nm Z22 was being added to SX each iteration, for a total of 400 nm Z22. Red estimates for Z22 were plus/minus and not consistently at this limit. What is wrong with SX/Blue?

02:45 While waiting for temps to stabilize... Testing operation of blue only... random field at 20h, 25deg. Focus/Collim... OK. Now blueonly.ob takes 5 dithers each 150sec... guiding and I see guide corrections being sent, but no Z2, Z3 in PSF GUI. So it is working as expected. MCS is running on 303. PCS on 302. PMC LEFT 301, PMC RIGHT 304, PSF LEFT 303, PSF RIGHT 303.

03:21 Slewed to NGC110. The goal of this test is to see whether Z4 corrections from tech chip are helping - effectively tracking the change in focus. We will runs OBs of duration 15min, 30min, 60min and collimate before/after. The Z4 correction needed after the OB should be minimal, if the tech chip active optics Z4 correction is working well.

03:23 Focus on NGC110 field. Clouds at first --- 033317, 033317 are the first pupils of dofpia.

-4057 -16 226 253 -106 247 51 1.11 1.61 B -2204 -79 256 387 -427 92 -27 0.98 0.91 R

03:39 Starting NGC110 sequence with 5 dithers, each 150 sec, for a total of 750 or 12.5 min --- about 15 min duration of OB. Between 1st and 2nd dither, I see that AO corrections are being sent: in lbc.log we see
2009/10/23 03:42:31.691942 S B TRACKERS          Active Optics P03 -418.32000000 +0.00000000
2009/10/23 03:42:36.487503 S R TRACKERS          Active Optics P03 -151.14000000 +0.00000000
and on PSF GUI LEFT and RIGHT I see the same Z4 values, -418.32 LEFT and -151.14 RIGHT. The LEFT value is wiped out soon, because guiding corrections are being sent to the SX mirror. The RIGHT value remains.

Next set, before 3rd dither:
2009/10/23 03:45:46.563337 S B TRACKERS          Active Optics P03 -520.50000000 +0.00000000
2009/10/23 03:45:47.353822 S R TRACKERS          Active Optics P03 +317.40000000 +0.00000000

Some clouds and fuzzier seeing during 03:49:30 image, so no focus corrections for that one, but the following, before 5th is good:
2009/10/23 03:48:57.108567 S R TRACKERS          Active Optics P03 -106.63000000 +0.00000000
2009/10/23 03:48:57.774068 S B TRACKERS          Active Optics P03 +288.98000000 +0.00000000
2009/10/23 03:52:03.993325 S R TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2009/10/23 03:52:05.997520 S B TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2009/10/23 03:55:10.644918 S R TRACKERS          Active Optics P03 -52.41000000 +0.00000000
2009/10/23 03:55:13.158654 S B TRACKERS          Active Optics P03 +587.57000000 +0.00000000

03:58: dofpia to Focus/collimate once that ended.

-724 -164 54 157 67 6 -27 0.90 0.31 R -867 49 296 289 84 -76 -98 1.05 0.39 B

  1. 114 -345 -94 22 -80 -15 1.15 0.16 B blue was OK, but R was messed up ... data took too long to appear and
I hit before file did arrive, so the result was that it sent correction 2x.

results from lbcfpiaBlueLog and RedLog:
 -1140      0      0      0      0    600      0    y  0.00  0.00                     dofpia_sph3_adjustment             
  -867     49    296    289     84    -76    -98    y  1.05  0.39                      /newdata/lbcb.20091023.035716.fits  V-BESSEL
   171    114   -345    -94     22    -80    -15    y  1.15  0.16                      /newdata/lbcb.20091023.035934.fits  V-BESSEL
   517     81   -286   -117    -11    -82     75    y  1.37  0.25                      /newdata/lbcb.20091023.040222.fits  V-BESSEL
  -990    128    313     32    -44    246    -14    y  1.17  0.44                      /newdata/lbcb.20091023.040407.fits  V-BESSEL
  -211   -127   -215    197     75    -92    -52    y  1.12  0.18                      /newdata/lbcb.20091023.040603.fits  V-BESSEL
  1312      0      0      0      0   -690      0    y  0.00  0.00                     dofpia_sph3_adjustment             
  -950      0      0      0      0    500      0    y  0.00  0.00                     dofpia_sph3_adjustment             
  -724   -164     54    157     67      6    -27    y  0.90  0.31                      /newdata/lbcr.20091023.035712.fits  r-SLOAN 
  -724   -164     54    157     67      6    -27    y  0.90  0.31                      /newdata/lbcr.20091023.035712.fits  r-SLOAN 
  1218    556   -288   -182    -89   -113      8    y  1.20  0.53                      /newdata/lbcr.20091023.040218.fits  r-SLOAN 
  -564    -92    168      9     32     43      4    y  1.02  0.24                      /newdata/lbcr.20091023.040405.fits  r-SLOAN 
  -128   -201      1     11     31     72    -11    y  0.97  0.12                      /newdata/lbcr.20091023.040600.fits  r-SLOAN 
   784      0      0      0      0   -412      0    y  0.00  0.00                     dofpia_sph3_adjustment             

04:15 Ran dofpia one more time. Now sending 30-min long OB on NGC110. Temperature differentials are: Glass therms = 2.462, Ambient therms 2.412 for Right and Glass=2.795, ambient = 2.619 for Left. WS ambient = 1.9
2009/10/23 04:18:18.460424 S R TRACKERS          Active Optics P03 +23.47000000 +0.00000000
2009/10/23 04:18:20.474853 S B TRACKERS          Active Optics P03 -363.10000000 +0.00000000
2009/10/23 04:21:29.478105 S R TRACKERS          Active Optics P03 +577.92000000 +0.00000000
2009/10/23 04:21:32.244710 S B TRACKERS          Active Optics P03 +48.87000000 +0.00000000
2009/10/23 04:24:42.238382 S R TRACKERS          Active Optics P03 -300.68000000 +0.00000000
2009/10/23 04:24:44.314498 S B TRACKERS          Active Optics P03 +522.23000000 +0.00000000
2009/10/23 04:27:51.377401 S R TRACKERS          Active Optics P03 +80.25000000 +0.00000000
2009/10/23 04:27:54.842353 S B TRACKERS          Active Optics P03 -587.24000000 +0.00000000
2009/10/23 04:31:04.614922 S R TRACKERS          Active Optics P03 +286.61000000 +0.00000000
2009/10/23 04:31:09.547725 S B TRACKERS          Active Optics P03 -674.41000000 +0.00000000
2009/10/23 04:34:18.739301 S R TRACKERS          Active Optics P03 -265.84000000 +0.00000000
2009/10/23 04:34:23.276995 S B TRACKERS          Active Optics P03 +505.02000000 +0.00000000
2009/10/23 04:37:30.754483 S R TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2009/10/23 04:37:35.016135 S B TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2009/10/23 04:40:37.158731 S R TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2009/10/23 04:40:40.884368 S B TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2009/10/23 04:43:44.323582 S R TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2009/10/23 04:43:46.941959 S B TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2009/10/23 04:46:48.758864 S R TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2009/10/23 04:46:52.866052 S B TRACKERS          Active Optics P03 +0.00000000 +0.00000000

Cloud again around 043459 - guiding failed for both blue and red.

04:50 OB finished - Ran FPIA but no pupil found. Will repeat at new field that is to the S - Clouds are coming from NW, but right now appear to cover all sky.

05:00 Went to new field in South. Ran dofpia.
 -1140      0      0      0      0    600      0    y  0.00  0.00                     dofpia_sph3_adjustment             
  1158    617    539   -200     34   -178   -100    y  1.01  0.54                      /newdata/lbcb.20091023.045900.fits  V-BESSEL
   541    193     65     28    -45   -195    -25    y  1.09  0.27                      /newdata/lbcb.20091023.050046.fits  V-BESSEL
  -162     -2   -361   -120    -44     15      8    y  1.13  0.16                      /newdata/lbcb.20091023.050241.fits  V-BESSEL
  1319      0      0      0      0   -694      0    y  0.00  0.00                     dofpia_sph3_adjustment             

  -950      0      0      0      0    500      0    y  0.00  0.00                     dofpia_sph3_adjustment             
  -594   -397    -93   -633    -94   -158    -53    y  0.82  0.41                      /newdata/lbcr.20091023.045901.fits  r-SLOAN 
   556    176    -87    464    111   -255    -12    y  0.81  0.36                      /newdata/lbcr.20091023.050043.fits  r-SLOAN 
   111    113     55    -83   -127   -174      6    y  0.88  0.16                      /newdata/lbcr.20091023.050238.fits  r-SLOAN 
   727      0      0      0      0   -383      0    y  0.00  0.00                     dofpia_sph3_adjustment             

Now running 30-min long OB. DDO221_30min. Blue Guide Star is galaxy for 1st dither position. So far it has only made a correction to Z4 once - and on Red only. Not enough stars on tech chip, esp with clouds, I think.
2009/10/23 05:07:56.998795 S R TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2009/10/23 05:07:57.242987 S B TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2009/10/23 05:11:00.004409 S R TRACKERS          Active Optics P03 +665.28000000 +0.00000000
2009/10/23 05:11:06.185833 S B TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2009/10/23 05:14:09.256404 S R TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2009/10/23 05:14:13.706157 S B TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2009/10/23 05:17:17.587679 S R TRACKERS          Active Optics P03 -633.18000000 +0.00000000
2009/10/23 05:17:21.302337 S B TRACKERS          Active Optics P03 -384.80000000 +0.00000000
2009/10/23 05:20:29.689420 S R TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2009/10/23 05:20:32.081778 S B TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2009/10/23 05:23:34.601640 S R TRACKERS          Active Optics P03 +636.38000000 +0.00000000
2009/10/23 05:23:39.697918 S B TRACKERS          Active Optics P03 -763.50000000 +0.00000000
2009/10/23 05:26:47.664830 S R TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2009/10/23 05:26:51.382400 S B TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2009/10/23 05:29:54.210581 S R TRACKERS          Active Optics P03 +234.77000000 +0.00000000
2009/10/23 05:29:58.639155 S B TRACKERS          Active Optics P03 +606.53000000 +0.00000000
2009/10/23 05:33:02.993131 S R TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2009/10/23 05:33:06.442247 S B TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2009/10/23 05:36:09.972534 S R TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2009/10/23 05:36:12.600170 S B TRACKERS          Active Optics P03 -564.53000000 +0.00000000

Ran dofpia after the sequence - took 2 iterations to converge for B and R.
  -950      0      0      0      0    500      0    y  0.00  0.00                     dofpia_sph3_adjustment             
  -631    407   -415   -132    180   -374     16    y  0.99  0.40                      /newdata/lbcr.20091023.053819.fits  r-SLOAN 
  -108    -47    -94    114     70   -320     46    y  0.89  0.23                      /newdata/lbcr.20091023.054030.fits  r-SLOAN 
   734      0      0      0      0   -386      0    y  0.00  0.00                     dofpia_sph3_adjustment             
 
-1140      0      0      0      0    600      0    y  0.00  0.00                     dofpia_sph3_adjustment             
  1741   -383   -655    -22   -197   -229     44    y  1.22  0.75                      /newdata/lbcb.20091023.053822.fits  V-BESSEL
  -372     70    239    -94    148     51     17    y  1.15  0.20                      /newdata/lbcb.20091023.054033.fits  V-BESSEL
  1337      0      0      0      0   -703      0    y  0.00  0.00                     dofpia_sph3_adjustment             


22:45 dofpia on new field that has more stars - so more likely to get a star on both Guide & ActOpt chips. NGC1023.
 -1140      0      0      0      0    600      0    y  0.00  0.00                     dofpia_sph3_adjustment             
 -1209   -377    639    356   -295    176     44    y  1.31  0.58                      /newdata/lbcb.20091023.054732.fits  V-BESSEL
  -917   -781    -21   -115    -42     61    -76    y  1.19  0.44                      /newdata/lbcb.20091023.054859.fits  V-BESSEL
   -96    241    349    -93     70    175    -41    y  1.15  0.20                      /newdata/lbcb.20091023.055053.fits  V-BESSEL
  1335      0      0      0      0   -702      0    y  0.00  0.00                     dofpia_sph3_adjustment             

  -950      0      0      0      0    500      0    y  0.00  0.00                     dofpia_sph3_adjustment             
 -1030   -326    553     93   -356    107     -5    y  1.00  0.49                      /newdata/lbcr.20091023.054732.fits  r-SLOAN 
  -991    -64    299   -150     60    431    -43    y  0.74  0.50                      /newdata/lbcr.20091023.054900.fits  r-SLOAN 
    87    246   -260    318    -40     53    -22    y  0.75  0.20                      /newdata/lbcr.20091023.055049.fits  r-SLOAN 
   644      0      0      0      0   -339      0    y  0.00  0.00                     dofpia_sph3_adjustment             

Took a while to converge - Blue pupils went through odd shapes (astig).

22:54 Starting NGC1023 OB to test once more focus tracking during a 30-min OB - this time one with more stars. There are pupils on the tech chips, but odd-shaped, not round. But, this time no corrections are being sent.

2009/10/23 05:56:05.115035 S R TRACKERS          Active Optics P03 +507.70000000 +0.00000000
2009/10/23 05:56:10.824794 S B TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2009/10/23 05:59:14.411673 S R TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2009/10/23 05:59:18.398160 S B TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2009/10/23 06:02:23.466330 S R TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2009/10/23 06:02:27.207416 S B TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2009/10/23 06:05:31.026299 S R TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2009/10/23 06:05:35.230286 S B TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2009/10/23 06:08:39.894607 S R TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2009/10/23 06:08:43.745014 S B TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2009/10/23 06:11:46.470814 S R TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2009/10/23 06:11:52.230912 S B TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2009/10/23 06:14:56.845286 S R TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2009/10/23 06:15:00.566096 S B TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2009/10/23 06:18:03.840001 S R TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2009/10/23 06:18:09.309835 S B TRACKERS          Active Optics P03 -540.50000000 +0.00000000
2009/10/23 06:21:18.987634 S R TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2009/10/23 06:21:22.022206 S B TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2009/10/23 06:24:25.978589 S R TRACKERS          Active Optics P03 +0.00000000 +0.00000000
2009/10/23 06:24:29.727458 S B TRACKERS          Active Optics P03 -528.69000000 +0.00000000

06:25 OB finished - running dofpia now.
 -1140      0      0      0      0    600      0    y  0.00  0.00                     dofpia_sph3_adjustment             
   789    378    -53   -141     68    -78      1    y  1.13  0.35                      /newdata/lbcb.20091023.062556.fits  V-BESSEL
   492   -283   -249    -81      4    -36     -8    y  0.93  0.24                      /newdata/lbcb.20091023.062822.fits  V-BESSEL
   376    235     96     65    -81   -166      1    y  1.02  0.22                      /newdata/lbcb.20091023.063046.fits  V-BESSEL
  1245      0      0      0      0   -655      0    y  0.00  0.00                     dofpia_sph3_adjustment             

  -950      0      0      0      0    500      0    y  0.00  0.00                     dofpia_sph3_adjustment             
  2360    -83   -196    275    101  -1000     11    y  0.66  1.14                      /newdata/lbcr.20091023.062556.fits  r-SLOAN 
   338   -146    224   -207    111   -187      0    y  0.66  0.24                      /newdata/lbcr.20091023.062819.fits  r-SLOAN 
   850   -217   -374     88     56   -441      6    y  0.71  0.46                      /newdata/lbcr.20091023.063043.fits  r-SLOAN 
   617      0      0      0      0   -325      0    y  0.00  0.00                     dofpia_sph3_adjustment             


Stopping this experiment. The last OB did put stars on the tech chips, but many times the corrections were not sent.

23:45 New experiment. Step through Z11, from -2000 to +2000 in steps of 500nm. Take extra/intra focal pupils at each value of Z11. Slew to open cluster in And 01:58:00, 37:41.

Focus/collimate:

Send -2000nm Z11 to SX and DX: run RB_rVexin4 --- went into Z limit on DX. At elev = 83, going to lower elevation.

00:00 New field - running dofpia on it.

Red controller error:

2009/10/23 07:09:52.163000 E R CAMERA CCDCTRL "CCD Controller fiber link error or incomplete data" (code:22) controller error [ccdctrl.c:2378] 2009/10/23 07:09:52.179000 E R CAMERA CCDCTRL "DLL allocation failed" (code:8) controller error [ccdctrl.c:1839]

Tried to re-run OB and it works... no need this time to power cycle camera.

Dofpia:
 -950      0      0      0      0    500      0    y  0.00  0.00                     dofpia_sph3_adjustment             
   783    364     33   -166     31     11    -16    y  0.62  0.34                      /newdata/lbcr.20091023.071432.fits  r-SLOAN 
   -72   -111      1     25     64     -3     -7    y  0.53  0.09                      /newdata/lbcr.20091023.071705.fits  r-SLOAN 
   497      0      0      0      0   -262      0    y  0.00  0.00                     dofpia_sph3_adjustment             

 -1140      0      0      0      0    600      0    y  0.00  0.00                     dofpia_sph3_adjustment             
  1399    697   -199    412    155   -560    -33    y  0.93  0.72                      /newdata/lbcb.20091023.071436.fits  V-BESSEL
   288    279    -99     -3    -97   -152      0    y  0.87  0.19                      /newdata/lbcb.20091023.071708.fits  V-BESSEL
  1145      0      0      0      0   -602      0    y  0.00  0.00                     dofpia_sph3_adjustment             

07:20: -2000 to both SX and DX. Ran RB_rVexin4.ob

B 072029 extrafoc (-0.8mm)

B 072118 intrafoc (+0.8mm)

R 072026 extrafoc

R 072116 intrafoc

removed -2000 (by sending +2000), dofpia again

07:25 +2000 to both SX and DX. RB_rVexin4.ob

B 073022 extrafoc

B 073111 intrafoc

R 073019 extrafoc

R 073111 intrafoc

Give -2000 nm Z11 to both sides to return to 0, and run dofpia again.

07:38: -500 to both SX and DX. RB_rVexin4.ob

B 073854 extrafoc

B 073943 intrafoc

R 073844 extrafoc

R 073934 intrafoc

Give +500nm to return to 0 and run dofpia.

07:45 +500 to both SX and DX. RB_rVexin4.ob

B 074501 extrafoc

B 074551 intrafoc

R 074458 extrafoc

R 074545 intrafoc

Give -500nm to return to 0 and run dofpia.

07:51 -1500 Z11 to SX and DX. RB_rVexin4.ob

B 075224 extrafoc

B 075313 intrafoc

R 075221 extrafoc

R 075310 intrafoc

Give +1500nm Z11 to return to 0 and run dofpia.

07:57 +1500 Z11 to SX and DX. RB_rVexin4.ob

B 075911 extrafoc

B 080001 intrafoc

R 075909 extrafoc

R 080000 intrafoc

Give -1500nm Z11 to return to 0 and run dofpia

08:07 -1000 Z11 to SX and DX. RB_rVexin4.ob

B 080810 extrafoc

B 080901 intrafoc

R 080807 extrafoc

R 080900 intrafoc

Give +1000m Z11 to return to 0 and run dofpia...

(note: 080452 on blue side has lots of Z22 evident).

08:12 +1000 Z11 to SX and DX. RB_rVexin4.ob

B 081457 extra

B 081547 intra

R 081454 extra

R 081545 intra

Give -1000nm Z11 to return to 0 and run dofpia.

08:24 0 Z11 RB_rVexin4.ob

B 082418 extra

B 082508 intra

R 082413 extra

R 082502 intra

08:25 Clearing forces on SX... Andrew gave -100nm Z22 to SX, then -50nm Z22 to SX.

08:27 Taking another RB_rVexin4.ob (compare with previous set).

B 082751 extra

B 082839 intra

R 082744 extra

R 082833 intra

08:31 Running dofpia now --- see if it tries to remove Z22 that Andrew entered. It did give +75 nm Z22... so the Blue Z22 zero point is incorrect and needs adjustment. Running dofpia again... does it still want to send more positive Z22? yes, +100nm. Total forces on Z22 are high -

08:43 -400 Z22 on SX and DX... RB_rVexin4

B 084452 extra

B 084543 intra

R 084452 extra

R 084543 intra

Giving +400 Z22 to return to 0 and running dofpia...

08:51 +400 Z22 on SX and DX... RB_rVexin4

B 085331 extra

B 085420 intra

R 085328 extra

R 085417 intra

Giving -400 z22 to return to zero and running dofpia...

08:59 -100 Z22 on SX and DX. RB_rVexin4

B 090016 extra

B 090105 intra

R 090011 extra

R 090103 intra

Giving +100 to return to zero, dofpia...

09:06 +100 Z22 on SX and DX RB_rVexin4.

B 090729 extra

B 090819 intra

R 090725 extra

B 090816 intra

Giving -100 Z22 to return to zero and dofpia...

09:13 -300 Z22 to SX and DX. RB_rVexin4

B 091403 extra

B 091453 intra

R 091400 extra

R 091449 intra

Giving +300 Z22 to return to 0 and running dofpia.

09:21 +300 Z22 to both SX and DX. RB_rVexin4.

B 092209 extra

B 092259 intra

R 092206 extra

R 092256 intra

Giving -300nm and running dofpia...

09:33 -200 Z22 to both SX and DX. RB_rVexin4

B 093354 extra

B 093442 intra

R 093350 extra

R 093440 intra

Giving +200 to return to 0 and running dofpia... elevation ~50 deg and taking a while to converge.

09:43 +200 Z22 to both SX and DX. RB_rVexin4

B 094356 extra

B 094447 intra

R 094353 extra

R 094444 intra

Giving -200 nm to return to zero.

Going to a new field, since this one is low.

09:52 Slewing to TU Tau. Going to take some extra/intra focal sequences with pupils of diameter 100pxls. RB_rV100exin4.ob Running dofpia.

* after collimation: RB_rV100exin4:

B 100118 extra (104 pixl)

B 100222 intra

R 100115 extra

R 100219 intra

Dofpia again..

10:06 -1000 Z4 to both SX and DX. RB_rV100exin4.

B 101155 extra

B 101259 intra

R 101147 extra

R 101253 intra

Add +1000 Z4 to go back to zero, dofpia again...

10:20 +1000 Z4 to both SX and DX. RB_rV100exin4.

B 102038 extra

B 102143 intra

R 102035 extra

R 102136 intra

giving -1000 Z4 and running dofpia again...

10:29 -1000 Z11 to both SX and DX. RB_rV100exin4.

B 103123 extra

B 103227 intra

R 103120 extra

R 103231 intra

giving +1000nm Z11 and running dofpia again...

10:41 +1000nm Z11 to both SX and DX. RB_rV100exin4.

B 104207 extra

B 104311 intra

R 104202 extra

R 104304 intra

give -1000nm Z11 and running dofpia again...

10:50 -300 Z22 to both SX and DX. RB_rV100exin4.

B 105104 extra

B 105210 intra

R 105101 extra

R 105213 intra --- right main mirror failed to focus and started exp anyway (problem with hp 3 --- not collim for exposure).

giving +300 Z22 and running dofpia...

11:00 +300 Z22 to both SX and DX. RB_rV100exin4.

B 110101 extra

B 110204 intra

R 110058 extra

R 110200 intra

giving -300 Z22 and running dofpia...

11:10 -1000 Z5 to both and RB_rV100exin4.

B 111120 extra

B 111223 intra

R 111117 extra

R 111224 intra

giving +1000 Z5 and running dofpia again...

11:20 +1000 Z5 to both and RB_rV100exin4

B 112237 extra

B 112339 intra

R 112234 extra

R 112336 intra

giving -1000 Z5 and running dofpia...

11:30 -1000 Z7 to both SX and DX, RB_rV100exin4.

B 113200 extra

B 113303 intra

R 113155 extra

R 113257 intra

giving +1000 Z7 and dofpia...seeing worsening --- 1.25 Blue and ~1 Red estimated by FPIA... well, no. it is just taking longer to converge. Maybe seeing is variable because next iteration estimates were 0.8 Blue and 0.6 Red

11:43 +1000 Z7 to both SX and DX. RB_rV100exin4.

B 114411 extra

B 114513 intra

R 114408 extra

R 114511 intra

gave -1000 nm z7 and run dofpia again...

11:48 done with taking extra/intra focal 100 pixel pairs. going now to repeat +/- 1000 Z5 and Z7 for the 60 pixel pairs.

11:50 -1000 Z5 RB_rVexin4.ob

B 115328 extra

B 115417 intra

R 115325 extra

R 115414 intra

giving +1000 Z5 and running dofpia...

12:00 +1000 Z5 RB_rVexin4.ob

B 120137 extra

B 120225 intra

R 120132 extra

R 120223 intra

giving -1000 Z5 and running dofpia...

12:08 -1000 Z7 to SX and DX. RB_rVexin4.

B 120913 extra

B 121003 intra

R 120911 extra

R 121002 intra

giving +1000 Z7 and running dofpia...

12:16 +1000 Z7 to SX and DX. RB_rVexin4.

B 121716 extra

B 121804 intra

R 121713 extra

R 121802 intra

gave -1000 Z7 and ran dofpia

12:20 Focus/collimate: new experiment to try to replicate the occassional vibrations in an attempt to analyze how these might affect FPIA.

* 122619, 122616 are extrafoc pupils taken while Wayne nudged telescope by 1" in declination. Do not see major change in astigmatism due to movement of telescope.

12:31 Second try, this time with nudge of 2"in the middle 123228(B) and 123225(R).

732 -959 1083 -33 -166 0 0 y 2.06 0.53 /newdata/lbcb.20091023.123228.fits V-BESSEL

  1. -313 399 -45 39 25 70 y 1.46 0.28 /newdata/lbcr.20091023.123225.fits r-SLOAN

12:38 another try again with 1" only: 123837 (R) and 123830(B)

-186 159 261 -128 -20 -27 1 y 1.17 0.15 /newdata/lbcr.20091023.123827.fits r-SLOAN -132 -240 511 -20 17 -15 57 y 1.37 0.19 /newdata/lbcb.20091023.123830.fits V-BESSEL

We need to analyze these pupils to see if there is an intensity difference in the major and minor axis.

05:49 sky flats at g-SLOAN & r-SLOAN. Sky brightness 15.00 when OB started - test image gave about 10k counts. Pointing at meridian and elevation 80 degrees. 125942 (B), 125939(R)... exposure times 1, then scaling by 0.5, 0.3 and 0.1.

13:11 Closing.


NOTES:
  • Blue Z22 zero point needs to be changed. An estimate is that it should be the current value minus 170nm
  • FPIA logs in /home/LBTO/ and are lbcfpiaBlueLog and lbcpfiaRedLog

-- OlgaKuhn - 23 Oct 2009
Topic revision: r3 - 23 Oct 2009, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback